### v5

@@ 4 Computational experiments
In this section, some computational experiments are presented. They consist of a set of random instances that test Algorithm 1 against attempting to solve IVPRM exactly, and the resolution of a real-world instance is addressed at the end.
The MIP formulations of the IVPRM and the R-IVPRM have been implemented in Gurobi 9.0.2 using the Python Programming Language. The tests were performed on a MacBook Air 2018 with Core i5 (1.6GHz) and 8GB of memory.
@@ 4.1 Simulated instances
Random instances were generated using information from the OpenStreetMaps API with the package osmnx [44]. Several Points of Interest (e.g.. entertainment venues and shops) were sampled from an urban neighborhood, covering a $1000$ m radius from its center. For each pair of stores, the shortest vehicular and pedestrian paths between them were computed. Then, the average travel time between stores was computed using an average speed of $30$ km/h for vehicles and $5$ km/h for pollsters. As vehicles must respect the sense of the street network, the resulting vehicle travel time matrix is asymmetric. Store-polling times, in minutes, were also randomly generated from a uniform distribution $\mathcal{U}(1,20)$.
As presented in Table 3, 10 instances were simulated with different values of $n$, $B_{\max}$ in minutes, $P$ in minutes, and $Q$. The number of available vehicles, pollsters, and days to solve each instance was computed using the nonlinear model (7a)–(7e). This was done by initializing $|K|=3$, $|E|=5$, and $|S|=15$, coinciding with the available real-world resources. Finally, the costs for the objective function were set to $\kappa_{0}=\$200$, $\kappa_{1}=\$100$, and $\kappa_{2}=\$40$.
@@
00footnotetext: Here, $|K|$ was bounded by $3$, $|E|$ by $5$, and $|S|$ by $15$.
The results of the experiments are summarized in Tables 4 and 5. In the former, different strategies for solving the IVPRM for each instance are reported alongside the bounds obtained following the procedures presented in Sections 2.5 and 2.4. For an aggregated daily plan, each strategy presents the number of vehicles used $\hat{k}$, the number of active pollsters $\hat{e}$, the number of days in the plan $\hat{s}$, the cost of the plan, and the optimality gap. The first three quantities are computed as follows:
$\hat{s}$ $:=\sum_{s\in S}u^{s}$ (11a)
$\hat{k}$ $:=\sum_{s\in S}\sum_{k\in K}\sum_{j\in C}y_{0,j}^{k,s}$ (11b)
$\hat{e}$ $:=\sum_{s\in S}\sum_{e\in E}\sum_{j\in C}z_{0,j}^{e,s}$ (11c)
@@
$\textrm{GAP}=\frac{|\mathtt{Objective}-\mathtt{Low}|}{\mathtt{Objective}},$
where $\mathtt{Objective}$ is the value of (1) and $\mathtt{Low}$ is the lower bound obtained from the branch and bound process.
The first strategy, MIP Solve, solves each instance of IVPRM within a runtime of $3000$ s. The second strategy, Heuristic Solve, applies Algorithm 1 as a solution method with a runtime of $1000$ s for each instance of R-IVPRM. The last strategy, Fixing Variables Heuristic, works as follows: the best feasible or optimal solution of R-IVPRM is retrieved for some initial partition, and the integer variables are fixed as an input for the IVPRM omitting the variables with ending nodes with $\{0,2n+1\}$. Then, IVPRM is attempted to be solved with those fixed variables to find an integer feasible solution in a bounded time. The procedure continues with the second partition if a solution is not found. It continues fixing and adding variables to the model until the best feasible or the optimal solution is obtained in a bounded time. It is important to remark that the Fixing Variable Heuristic produces a marginal improvement compared to the second strategy. Only for instances with $n=14$ and $n=16$, the latter strategy outperformed Algorithm 1. Notice that $\mathtt{Low}$ coincides with the lower bound obtained from Section 2.5. Moreover, MIP Solve and Fixing Variables Heuristic were not able to find any improved lower bounds.
@@
$n$ Bounds MIP Solve Heuristic Solve Fixing Variables
@@
00footnotetext: Each row represents the results for an instance of size $n$ described in Table 3.
A detailed report of the second strategy is presented in Table 5. Each set of stores was partitioned into $2$ or $4$ subsets, labeled A to D. For each subset, the number of vehicles used $\bar{k}$, the number of active pollsters $\bar{e}$, the cost of the half shift, and the optimality gap are included. The first two quantities are computed as $\bar{k}:=\sum_{k\in K}\sum_{j\in C}y_{0,j}^{k}$ and $\bar{e}:=\sum_{e\in E}\sum_{j\in C}z_{0,j}^{e}$. Only the partitions where R-IVPRM had a feasible solution for all subsets are included. A final column is included with the cummulative solution times (in seconds).
Table 5: Results of R-IVPRM for simulated instances at each subset. Each row represents the results for an instance of size $n$ described in Table 3.
@@
00footnotetext: Each row represents the results for an instance of size $n$ described in Table 3.
@@ 4.2 A real-world case study
The case study proposed by INEC is focused on the city of Guayaquil. As the country’s main commercial port, its relevancy in economic and populational development makes this city of 3 million inhabitants ideal for collecting data to monitor the CPI. Hence, INEC deploys a group of at most $5$ pollsters and $3$ vehicles monthly to collect data from $820$ stores in Guayaquil and its outskirts, see Figure 4. Currently, the planning is done empirically using all the available resources and is completed in $17$ days. The objective of the operational managers at INEC was to reduce the costs associated with the number of vehicles and pollsters within a time horizon of, at most, $15$ days. Furthermore, on each day, the workforce must finish its duties in at most $7$ hours with an included break of $40$ minutes. Accordingly, $n=820$, $S=\{0,\ldots,14\}$, $E=\{0,1,2,3,4\}$, $K=\{0,1,2\}$, $B_{\max}=420$ min, and $P=40$ min. Finally, $\kappa_{0}=\$200$, $\kappa_{1}=\$100$, and $\kappa_{2}=\$40$.
The travel time between each pair of stores was computed using the shortest path in the road and walking networks provided by a geographic information system (GIS). In the walking network, the average walking velocity of pollsters is fixed at $5$ km/h, whereas the average vehicle displacement velocity in the city is fixed at $30$ km/h. In contrast to the simulated instances, a $40$ km/h velocity is fixed for the outskirts of the city. Service times for each store were measured and provided by INEC. Computation of the lower bounds, as in Section 2.5, for the resulting instance of the IVPRM, yields that any feasible aggregated daily plan requires at least $3$ days, $15$ pollsters, and $4$ vehicles.
Figure 4: Map of Guayaquil displaying the stores to be visited.
@@ 4.3 Results and solution visualization
Upon the first attempt to solve the IP implementation of IVPRM for a small number of nodes in the graph ($n=15$), no feasible solution was found, even when using an HPC server. A similar result was obtained when attempting to use the variable fixing heuristic. These facts motivated the construction and application of the three-phase solution approach captured by Algorithm 1. The following tables describe the behavior of the latter algorithm using real-world data. It is important to remark that a lower bound for the number of partitions is considered as the ratio between the total service time and the total working time of the available pollsters per day ($\underline{k}=10$).
@@
Table 6 contains general information about partitions of the set of vertices of the graph for several values of $k$, which is the number of subsets in each partition. For each $k$ and $N_{\max}=5000$, the minimum number of vertices $n_{\min}$, the maximum number of vertices $n_{\max}$, and the minimum and maximum service times over all subsets in the partition are reported. The last column reports the number of subsets in the partition for which a feasible solution was found by the solver Gurobi in at most $T_{\max}=1$ h. Observe that for $k\in\{10,12,\dots,28\}$, Algorithm 1 can not construct a feasible daily plan since there are not as many solutions as subsets per partition.
Even though the maximum available number of pollsters and vehicles at INEC are $|E|=5$ and $|K|=3$, Algorithm 1 always reported that two pollsters and one vehicle are enough to construct every half-shift of a feasible daily plan. Therefore, Table 7 reports the time of the last vehicular arrival at the depot for each subset of the reported partition with $k=30$. For instance, the workload of subset $24$ is made in at most $2$ hours. It was noted that half of the instances were solved to optimality in less than two minutes, while the rest took the maximum allowed time.
@@
The results of the linking partition step are reported in Table 8. A pair of subsets are assigned to each day $s\in\{0,\ldots,14\}$, corresponding to a morning and afternoon half-shift needed to accomplish the workload. For example, day $11$ syncs subsets $27$ and $1$ for a total daily working time of $4$ h $53$ min, including a $40$ min break between shifts. The first subset of stores is polled in the morning, and the second is visited in the afternoon. As can be seen in Table 7, each daily plan is accomplished by $2$ pollsters and $1$ vehicle, concluding the workload in a time horizon of 15 days. The latter result contrasts with the empirical solution provided by INEC in which $5$ pollsters and $3$ vehicles were required in a time horizon of $17$ days. As the daily income of each pollster is \$$\$40$, and the daily cost of hiring and using a vehicle is \$$\$100$, the empirical solution currently used by INEC approximately duplicates the cost of the aggregated daily plan found with Algorithm 1.
@@ 7 $25\to 19$ 4 h 52 min Average: 5 h 03 min
Finally, the resulting real aggregated daily plan presented above is depicted in Figure 5. On the left upper corner, the afternoon shift of the eighth working day, subset 12, is zoomed in. The stores that the first and second pollsters visit are represented by circles and squares, respectively. Each store is labeled with a number between 1 and 820. Continuous and dashed lines represent routes, the former for pedestrians and the latter for vehicles.
@@ 5 Conclusions
The Integrated Vehicle and Pollster Routing Problem (IVPRP), focused on the transportation of pollsters with pedestrian and vehicular routes, has been introduced in this paper. It generalizes the Pickup and Delivery Problem, combining well-known problems such as the Share-a-Ride, the Integrated Dial-a-Ride, and the periodic VRP with time windows and backhaulings. The complexity of the IVPRP problem has been shown to be $\mathcal{NP}$–hard by reduction from the Traveling Salesman Problem.
A mixed integer formulation was proposed for this problem and assessed using real-world data. Due to the hardness of the problem, a three-phase heuristic method was devised: the set of stores to be visited is partitioned at the first stage, then pollsters and vehicles are routed on each subset, and finally, pairs of subsets are linked as half-shifts of an aggregated daily plan.
The algorithm was tested with a real-world instance at Guayaquil, Ecuador’s second most important city. INEC provided the data with information related to 820 stores to be visited monthly. The heuristically-found solution improved the one designed by INEC, reducing operational costs by approximately two times since half of the available workforce and fleet is required. Moreover, the resulting aggregated daily plan takes two days less than the former empirical solution. Despite the computational intractability of this problem, the previous results can be improved by working on tighter formulations and exact algorithms, which are interesting topics for future research.
@@
Acknowledgments The authors are grateful with the managers of INEC, specially with Alexandra Enríquez for providing the data used in the real-world instance, and for her valuable feedback on the results. Additionally, we acknowledge the collaboration of Pablo Zuleta during the first stage of this research.
@@
[…]
Data, code, and materials Code for the simulated instances and the implemented models is available at

### v6

@@ 4 Computational experiments
In this section, computational experiments are presented. These consist of a set of random instances that test Algorithm 1 against attempting to solve IVPRM exactly. Additionally, the resolution of a real-world instance is addressed at the end of the section.
The MIP formulations of IVPRM and R-IVPRM have been implemented in Gurobi 12 [ARM] using the Python programming language. The tests were performed on a MacBook Pro with an M1 processor and 16GB of memory. All instances, as well as the source code of the implementation, can be downloaded from:
@@ 4.1 Simulated instances
Random instances were generated using data from the OpenStreetMaps API with the osmnx package [48]. Several Points of Interest (e.g., entertainment venues and shops) were sampled from an urban neighborhood, covering a $1000$ m radius from its center. For each pair of stores, the shortest vehicular and pedestrian paths between them were computed. The average travel time between stores was then calculated using an average speed of $30$ km/h for vehicles and $5$ km/h for pollsters. Since vehicles must respect the street network directions, the resulting vehicle travel time matrix is asymmetric. Store-polling times, in minutes, were also randomly generated from a uniform distribution $\mathcal{U}(1,20)$.
As presented in Table 3, 10 instances were simulated with different values of $n$, $B_{\max}$ in minutes, $P$ in minutes, and $Q$. The number of available vehicles, pollsters, and days to solve each instance was computed using the nonlinear model Equation 7aEquation 7e. This was done by initializing $|K|=3$, $|E|=5$, and $|S|=15$, corresponding to the available real-world resources. Finally, the costs for the objective function were set to $\kappa_{0}=\$200$, $\kappa_{1}=\$100$, and $\kappa_{2}=\$40$.
@@
Here, $|K|$ was bounded by $3$, $|E|$ by $5$, and $|S|$ by $15$.
Four different strategies are used to solve the IVPRP, each focusing on a distinct approach to the problem. The first strategy involves solving the Mixed-Integer Linear Program (MIP) IVPRM using Gurobi in its default configuration. This B&B-based method runs for up to 3 00030003\,0003 000 seconds and aims to find optimal solutions when possible. The second procedure combines the B&B approach with Gurobi’s NoRel heuristic. The NoRel heuristic is a general-purpose technique for MIPs that skips solving the LP relaxation and instead uses methods like rounding and presolve reductions to identify feasible solutions. It is run for $120$ seconds, attempting to generate initial solutions. The third approach is the problem-tailored heuristic described in Algorithm 1. This approach is specifically designed for the structure of IVPRM and is particularly effective for large-scale instances due to its partitioning step. The fourth and final strategy is the Fixing Variables Heuristic. This method combines elements of the exact approach and the tailored heuristic. It starts by solving a relaxed version of IVPRM to find the best feasible or optimal solution for an initial partition. The integer variables from this solution are then fixed as inputs for IVPRM, except for those corresponding to the ending nodes $\{0,2n+1\}$. With the fixed variables, the model is solved again within a limited runtime to search for an integer feasible solution. If no solution is found, the process continues with the next partition, iteratively fixing and adding variables until the best feasible or optimal solution is achieved. This strategy is most suitable for smaller instances as it involves solving IVPRM, which does not scale well. The results of these strategies are summarized in Tables 4 and 5.
Table 4: Results of applying MIP model & Gurobi NoRel Heuristic
$n$ Bounds IVPRM NoRel Heuristic
$(\ell_{k},\ell_{e},\ell_{s})$ Low Up $(\hat{k},\hat{e},\hat{s})$ Obj. GAP $(\hat{k},\hat{e},\hat{s})$ Obj. GAP
10 (2,2,1) 480 620 (2,2,1) 480 0.0% (2,2,1) 480 0.0%
12 (2,2,1) 480 960 (2,2,1) 480 0.0% (2,2,1) 480 0.0%
14 (2,2,1) 480 960 (3,3,2) 820 41.5% (3,3,2) 820 41.5%
16 (1,2,1) 380 1040 (2,4,2) 760 50.0%
18 (1,2,1) 380 1040 (3,4,2) 860 55.8%
20 (1,2,1) 380 660 (2,4,1) 560 32.1%
25 (1,2,1) 380 960 (2,4,2) 760 50.0%
30 (1,2,1) 380 960 (2,4,2) 760 50.0%
40 (1,3,1) 420 1120 (4,8,2) 1120 62.5%
50 (2,6,2) 840 1400
Each row represents the results for an instance of size $n$ described in Table 3.
Table 4 compares the behavior of the Gurobi solver under different configurations as mentioned above. The first column presents the number of nodes in the instances. The following three columns, grouped in the multicolumn Bounds, report the bounds obtained following the procedures described in Sections 2.4 and 2.5. The other two multicolumns, IVPRM and NoRel Heuristic, each consist of three columns displaying the triplet $(\hat{k},\hat{e},\hat{s})$, the objective value, and the gap of the solution. Here, the triplet $(\hat{k},\hat{e},\hat{s})$ represents the cummulative number of vehicles used, the number of active pollsters, and the number of days in the plan, respectively. These quantities are computed as follows:
$\hat{s}\coloneqq\sum_{s\in S}u^{s},\qquad\hat{k}\coloneqq\sum_{s\in S}\sum_{k\in K}\sum_{j\in C}y_{0,j}^{k,s},\qquad\hat{e}\coloneqq\sum_{s\in S}\sum_{e\in E}\sum_{j\in C}z_{0,j}^{e,s}.$
@@
$\mathrm{GAP}=\frac{|\mathtt{Objective}-\mathtt{Low}|}{\mathtt{Objective}},$
where $\mathtt{Objective}$ is the value of Equation 1, and $\mathtt{Low}$ is the maximum between the lower bound obtained from Section 2.5 and the lower bound obtained from the branch-and-bound process.
Table 5 evaluates the behavior of the two heuristic methods, namely, Algorithm Algorithm 1 and the Fixing Variables Heuristic. The instance information in this table coincides with that of Table 4.
@@
$n$ Bounds Algorithm 1 Fixing Variables
@@
Each row represents the results for an instance of size $n$ described in Table 3.
Interesting insights can be drawn from the results. First, it is noteworthy that the customized invocation of the NoRel heuristic contributed significantly to finding feasible solutions in all but one instance. However, when comparing the NoRel heuristic with Algorithm 1 and the Fixing Variables Heuristic, NoRel outperforms Algorithm 1 only in instance 14. For all other instances, the proposed algorithms achieve better solutions than NoRel. It is important to remark that the Fixing Variables Heuristic produces a marginal improvement when compared to Algorithm 1.
For completeness, a detailed report on the computational results of Algorithm 1 is provided in Table 6. Each set of stores is partitioned into $2$ or $4$ subsets, labeled A through D. For each subset, the table includes the number of vehicles used $\bar{k}$, the number of active pollsters $\bar{e}$, the cost of the half shift, and the optimality gap. The first two quantities are computed as
$\bar{k}\coloneqq\sum_{k\in K}\sum_{j\in C}y_{0,j}^{k}\qquad\text{and}\qquad\bar{e}\coloneqq\sum_{e\in E}\sum_{j\in C}z_{0,j}^{e}.$
Only partitions where R-IVPRM had a feasible solution for all subsets are included. A final column is provided with the cumulative solution times (in seconds).
Table 6: Results of R-IVPRM for simulated instances at each subset.
@@
Each row represents the results for an instance of size $n$ described in Table 3.
@@ 4.2 A real-world case study
The case study proposed by INEC focuses on the city of Guayaquil. As the country’s main commercial port, its relevance in economic and population development makes this city of 3 million inhabitants ideal for collecting data to monitor the CPI. Hence, INEC deploys a group of at most $5$ pollsters and $3$ vehicles monthly to collect data from $820$ stores in Guayaquil and its outskirts (see Figure 4). Currently, the planning is done empirically using all available resources and data collection is completed in $17$ days.
The objective of the operational managers at INEC was to reduce the costs associated with the number of vehicles and pollsters within a time horizon of at most $15$ days. Furthermore, each day, the workforce must complete its duties within $7$ hours, including a $40$-minute break. Accordingly, the baseline parameters and sets are $n=820$, $S=\{0,\ldots,14\}$, $E=\{0,1,2,3,4\}$, $K=\{0,1,2\}$, $B_{\max}=420$ minutes, and $P=40$ minutes. Additionally, the cost parameters are set to $\kappa_{0}=\$200$, $\kappa_{1}=\$100$, and $\kappa_{2}=\$40$.
The travel time between each pair of stores was computed using the shortest path in the road and walking networks provided by a geographic information system (GIS). In the walking network, the average walking velocity of pollsters is fixed at $5$ km/h, whereas the average vehicle displacement velocity in the city is fixed at $30$ km/h. In contrast to the simulated instances, a velocity of $40$ km/h is fixed for the outskirts of the city. Service times for each store were measured and provided by INEC.
Computation of the lower bounds, as described in Section 2.5, for the resulting instance of IVPRM yields that any feasible aggregated daily plan requires at least $3$ days, $15$ pollsters, and $4$ vehicles.
Figure 4: Map of Guayaquil showing the locations of stores to be visited.
@@ 4.3 Results and solution visualization
Upon the first attempt to solve the IP implementation of IVPRM for a small number of nodes in the graph ($n=15$), no feasible solution was found, even when using an HPC server. Similarly, the variable fixing heuristic also failed to produce a solution. This outcome is not surprising, as the full model contains nearly 400 million variables, resulting in an LP that exceeds available memory. These challenges motivated the development and application of the three-phase solution approach described Algorithm 1.
The following tables summarize the behavior of this algorithm using real-world data. Notably, a lower bound for the number of partitions is determined as the ratio between the total service time and the total working time available per pollster each day ($\underline{k}=10$).
@@
[…]
Even though the maximum number of pollsters and vehicles available at INEC are $|E|=5$ and $|K|=3$, Algorithm 1 consistently reported that two pollsters and one vehicle are sufficient to construct every half-shift of a feasible daily plan. Consequently, Table 8 presents the time of the last vehicular arrival at the depot for each subset in the reported partition with $k=30$. For instance, the workload of subset $24$ is completed in at most $2$ hours. It was observed that half of the instances were solved to optimality in less than two minutes, while the remaining instances required the maximum allowed time.
@@
The results of the linking partition step are presented in Table 9. A pair of subsets is assigned to each day $s\in\{0,\ldots,14\}$, corresponding to a morning and afternoon half-shift required to accomplish the workload. For example, day $11$ syncs subsets $27$ and $1$ for a total daily working time of $4$ hours and $53$ minutes, including a $40$-minute break between shifts. The first subset of stores is polled in the morning, and the second is visited in the afternoon.
As shown in Table 8, each daily plan is executed by $2$ pollsters and $1$ vehicle, completing the workload within a time horizon of $15$ days. This result contrasts with the empirical solution provided by INEC, which required $5$ pollsters and $3$ vehicles within a time horizon of $17$ days. Given that the daily income of each pollster is \$$\$40$ and the daily cost of hiring and using a vehicle is \$$\$100$, the empirical solution currently used by INEC approximately doubles the cost of the aggregated daily plan found with Algorithm 1.
@@
Finally, the resulting real aggregated daily plan is depicted in Figure 5. In the upper left corner, the afternoon shift of the eighth working day (subset 12) is zoomed in. The stores visited by the first and second pollsters are represented by circles and squares, respectively, with each store labeled by a number between $1$ and $820$. Continuous and dashed lines indicate the routes: continuous lines represent pedestrian routes, while dashed lines represent vehicle routes.
@@ 5 Conclusions
The Integrated Vehicle and Pollster Routing Problem (IVPRP), focused on the transportation of pollsters with pedestrian and vehicular routes, has been introduced in this paper. It generalizes the Pickup & Delivery Problem, combining well-known problems such as the Share-a-Ride, the Integrated Dial-a-Ride, and the periodic Vehicle Routing Problem with time windows and backhaulings. The complexity of the IVPRP has been shown to be $\mathcal{NP}$–hard by reduction from the Traveling Salesman Problem.
A mixed integer formulation was proposed for this problem and assessed using simulated instances and real-world data. Due to the hardness of the problem, a three-phase heuristic method was devised: the set of stores to be visited is partitioned at the first stage, then pollsters and vehicles are routed on each subset, and finally, pairs of subsets are linked as half-shifts of an aggregated daily plan.
The algorithm was tested with a real-world instance at Guayaquil, Ecuador’s second most important city. INEC provided the data with information related to 820 stores to be visited monthly. The heuristically-found solution improved the one designed by INEC, approximately reducing twice the operational costs since half of the available workforce and fleet is required. Moreover, the resulting aggregated daily plan takes two days less than the former empirical solution.
Finally, future research could concentrate on the development of more advanced heuristic and metaheuristic approaches to address larger instances of the IVPRP. This could involve hybrid methods that integrate exact and heuristic techniques to achieve a balance between solution quality and computational efficiency. Additionally, there is significant potential to explore extensions of the IVPRP incorporating more complex constraints, such as dynamic routing adjustments based on real-time data or multi-period planning that accounts for demand variability over time. While this study focuses on a specific case in Ecuador, future research could investigate how the proposed framework might be adapted to similar applications in other fields, including logistics, urban planning, and disaster response.
@@
Acknowledgments The authors are grateful with the managers of INEC, specially with Alexandra Enríquez for providing the data used to build the real-world instance, and for her valuable feedback on the results. Additionally, we acknowledge the collaboration of Pablo Zuleta during the first stage of this research. We greatly appreciate the anonymous referees’ suggestions, which improved significantly the clarity and impact of the findings of this paper.